Five batch jobs, A through E, arrive in alphabetical order at a computer at almost the same time (i.e. they all arrive before the next scheduling event occurs). They have estimated running times of 500, 300, 600, 700, and 200 ms, respectively. Using diagrams, show how the process scheduling algorithms below would schedule these jobs, and calculate the mean process turnaround time (assuming no process switching overhead):

  1. FCFS
  2. SPN
  3. RR with a quantum of 100ms
  4. RR with a quantum of 150ms
Answer #1

A) FCFS :- A Process Assiralline Blast Time/ comptron Time 1111 C1-AT 5000 500 B 800-800 1400 700 2100 21000 2100 2300 230-02BT 4002000 Seb 200 le c) RR with TQ=100 - Process AT CT TAT=CT-AT A 1800 1800-0 = 1800 B 1200 1200-0 = 1200 = 2100 2100 2100-

Any queries please comment

Please thumbsup for my effort

Thank you and all the best